OpenStudy (anonymous):

how to find the integration int (sinh^-1x)dx

OpenStudy (amistre64):

it equals ln(x+sqrt(x^2+1))

OpenStudy (amistre64):

thats the derivative...lol

OpenStudy (anonymous):

i need the procedure.......whatever be the ans...

OpenStudy (anonymous):

i have books where the ans is given but not the procedure......

OpenStudy (amistre64):

ack!!...sinh^-1x = ln(x+sqrt(x^2+1))

OpenStudy (amistre64):

integrate the ln function; is that possible?

OpenStudy (nowhereman):

There is a better way, substitute x = sinh(z). Then you can integrate by hand, because only terms of the form exp(z) and z exp(z) remain.

OpenStudy (nowhereman):

I get \[z\cosh z - \sinh z = \sqrt{1+x^2}\sinh^{-1} x - x\] can that be correct?

OpenStudy (nowhereman):

oh wait, you have to exchange sinh and cosh there ;-) but then it's correct. Yeah that substitution works :-)